Please help me solve this equation: 2=n-(2n+3). I need to find the value of n. The book has the answer in the back for reference but I keep coming up with the wrong answer. This is the solution I got:
2=n-(2n+3)
2=n-2n+3
2=1n+3
-3=1n-3
-1/1=1n/1 n=1

Found 3 solutions by longjonsilver, scoop, shorty33353:
Answer by longjonsilver(2297)

you have 2 glaring errors in your working.

solution:
2=n-(2n+3)
2=n-2n-3
2=-n-3
2+n = -3
n = -5

jon.

Answer by scoop(27)

2=n-(2n+3)=
2=n-2n-3(the negative in front of parenthesis, changes all signs)
then add 3 to both sides
2+3=n-2n-3+3
5=n-2n
5=-n
divide bothe sides by the negative
n=-5
If you replace the n with a negative 5 you wll see that this works!!

Answer by shorty33353(1)

2=n-(2n+3)
2=n-2n-3
2+3=n-2n
5=n-2n
5=-1n
5/-1
n=-5